# You Only Learn One Query: Learning Unified Human Query for Single-Stage Multi-Person Multi-Task Human-Centric Perception

### v2

@@ 5.3 Results on the OCHuman Dataset
To verify the performance of HQNet in challenging crowded scenarios, we compare it with recent works on OCHuman dataset [89], which is a crowded scene benchmark for human detection, segmentation, and pose estimation. We show that our model outperforms previous methods under the same ResNet50 backbone network by a large margin. For instance, it outperforms SBL by 9.6 keypoint AP and CondInst by 11.0 segmentation AP on test set. It even achieves superior performance than HrHRNet (40.3 vs 40.0) and LOGO-CAP (40.3 vs 39.0) even with a much smaller backbone (ResNet-50 vs. HRNet-w32). With a stronger backbone, i.e. ViT-L, our HQNet sets new state-of-the-art results on detection (35.8 AP), segmentation (38.8 AP), and pose estimation (45.6 AP).
@@ 5.4 Generalize to New HCP Tasks
Unseen-task (zero-shot) generalization. We evaluate the generalization ability of our approach through an unseen task evaluation, specifically multiple object tracking (MOT) on the PoseTrack21 dataset [14]. Our models are trained solely on the COCO-UniHuman image-based dataset without explicit tuning for MOT. We hypothesize that our learned human query embeddings, which encode instance-specific features of diverse granularity, can serve as strong cues for distinguishing different objects. We utilized DeepSORT [77] and used the learned Human Query as re-identification features for association. In Table 7, we compare our results with two state-of-the-art single-network MOT methods that were pretrained on COCO and fine-tuned on PoseTrack21. Despite not explicitly being trained for MOT, our HQNet (R-50) achieves highly competitive results (64.6 IDF1 and 51.1 MOTA). This demonstrates the generalization ability of our learned Human Query. HQNet-Det refers to HQNet trained solely on the detection task, and we observed that co-training on multiple HCP tasks improved the quality of the query embeddings (64.6 IDF1 vs. 62.4 IDF1). Furthermore, by employing a stronger ViT backbone, our approach achieves state-of-the-art performance in the “zero-shot” MOT evaluation.
@@ 5.6 More Analysis
Effect of HQ-Ins Matching. “w/o HQ-Ins Matching” means using detection loss only for bipartite matching [85]. “w/ HQ-Ins Matching” means comprehensively using detection, pose, and segmentation loss for bipartite matching. As shown in Fig. 4, with detection only matching [85], there may be some erroneous cases when one person’s pose is matched to another person. HQ-Ins Matching avoids such errors by comprehensively considering multiple tasks as a whole. More quantitative evaluation can be found in Sec. A3.2.
@@ 6 Conclusion
In this work, we present a unified solution towards single-stage multi-task human-centric perception, called HQNet. The core idea is to learn a unified query representation (Human Query) that encodes different features (local and global appearance features, coarse and fine-grained localization features) for each instance. To facilitate model training and evaluation, we introduce a large-scale benchmark, termed COCO-UniHuman benchmark, to unify different representative HCP tasks (including classification, detection, segmentation, and keypoint localization). We extensively compare our proposed method with several state-of-the-art task-specific and multi-task approaches, and show the effectiveness of our proposed method.
Limitations and future work. While our work focuses on RGB image based HCP tasks, tasks about video data or multi-modality data (e.g. IR and Depth) also hold significant potential. We encourage future research to explore more comprehensive multi-task human-centric perception.
Acknowledgement. This paper is partially supported by the National Key R&D Program of China No.2022ZD0161000 and the General Research Fund of Hong Kong No.17200622.
Appendix A1 COCO-UniHuman Dataset Statistics
@@
Appendix A2 COCO-UniHuman Dataset Annotation
@@
A2.1 Body-based vs face-based annotation strategies.
@@
A2.2 Two-stage vs one-stage annotation strategies.
@@
Appendix A3 More Experimental Analysis
@@
A3.1 General class models vs person-specific models
@@
A3.2 Effect of HumanQuery-Instance Matching
@@
A3.4 Attention Visualization
Appendix A4 More Implementation Details
A4.1 Loss Functions
@@
A4.2 Details about Training
@@
A4.3 Details about Inference
@@
Appendix A5 Details about Baseline Models
Details about human detection baselines. For human detection, we compare two baseline approaches, i.e. Faster-RCNN [60] and DINO [85]. Note that these general object detectors are originally trained to handle general 80 classes (marked with * in Table A3). For fair comparisons, we use MMDetection [8] to re-train and evaluate them on “Person” category using the default experimental setting. Note that MMDetection re-implementation can be a little bit better than the original implementation.
Details about human pose estimation baselines. For human pose estimation, we compare with several representative top-down methods (SBL [78], HRNet [67], Swin [50] and PRTR [38]), bottom-up approaches (HrHRNet [10], DEKR [18], and SWAHR [52]) and single-stage approaches (FCPose [53], InsPose [66], PETR [65] and CID [73]). Note that the results of Swin (Swin-L) and CID (R-50-FPN) are from MMPose [12], and other results are from their original papers. Top-down methods generally yield superior performance, but often rely on a separate human detector, incurring redundant computational costs. Specifically, SBL, HRNet and Swin use the same person detector provided by [78], which is a strong Faster-RCNN [60] based detector with detection AP $56.4$ for the “Person” category on the COCO’2017 val set. PRTR applies a DETR-based person detector for human detection, which achieves $50.2$ AP for the whole “Person” category on the COCO’2017 val set. While PRTR introduces an end-to-end variant (E2E-PRTR) optimizing detection and pose jointly, it lags behind separately trained top-down approaches. For pose estimation, the input resolution for SBL, HRNet, and Swin is set as $256\times 192$, while the input resolution for PRTR is $384\times 288$. Bottom-up methods learn instance-agnostic keypoints and then cluster them into corresponding individuals. HrHRNet, DEKR, and SWAHR adopt the strong HRNet-w32 [67] backbone network with an input resolution of $512\times 512$. Single-stage approaches directly predict human body keypoints in a single stage. FCPose, InsPose and PETR adopt R-50 [21] backbone network. The input images are resized to have their shorter sides being 800 and their longer sides less or equal to 1333. For CID, we report both the results of R-50-FPN and HRNet-w32 backbones. The input resolution of CID is $512\times 512$.
@@
Details about mesh estimation baseline. For human mesh estimation, we compare HQNet with state-of-the-art one-stage monocular method ROMP [68]. Following official settings, we train ROMP on COCO-UniHuman with ResNet50 backbone.
Appendix A6 Discussion about Unifying HCP Tasks
A6.1 General network architecture design
@@
Appendix A7 Discussion about Human Attribute Recognition

### v3

@@ 5.3 Results on the OCHuman Dataset
To verify the performance of HQNet in challenging crowded scenarios, we compare it with recent works on OCHuman dataset [95], which is a crowded scene benchmark for human detection, segmentation, and pose estimation in Table 4. We show that our model outperforms previous methods under the same ResNet50 backbone network by a large margin. For instance, it outperforms SBL by 9.6 keypoint AP and CondInst by 11.0 segmentation AP on test set. It even achieves superior performance than HrHRNet (40.3 vs 40.0) and LOGO-CAP (40.3 vs 39.0) even with a much smaller backbone (ResNet-50 vs. HRNet-w32). With a stronger backbone, i.e. ViT-L, our HQNet sets new state-of-the-art results on detection (35.8 AP), segmentation (38.8 AP), and pose estimation (45.6 AP).
@@ 5.4 Generalize to New HCP Tasks
Unseen-task generalization. We evaluate the generalization ability of our approach through an unseen task evaluation, specifically multiple object tracking (MOT) on the PoseTrack21 dataset [15]. Our models are trained solely on the COCO-UniHuman image-based dataset without explicit tuning for MOT. We hypothesize that our learned human query embeddings, which encode instance-specific features of diverse granularity, can serve as strong cues for distinguishing different objects. We utilized DeepSORT [81] and used the learned Human Query as re-identification features for association. In Table 7, we compare our results with two state-of-the-art single-network MOT methods that were pretrained on COCO and fine-tuned on PoseTrack21. Despite not explicitly being trained for MOT, our HQNet (R-50) achieves highly competitive results (64.6 IDF1 and 51.1 MOTA). This demonstrates the generalization ability of our learned Human Query. HQNet (D) and HQNet (D+S) refers to HQNet trained solely on the detection (D) and segmentation (S) tasks respectively, and we observed that co-training on multiple HCP tasks improved the quality of the query embeddings (64.6 IDF1 vs. 62.4 IDF1). Furthermore, by employing a stronger ViT backbone, our approach achieves state-of-the-art performance.
@@ 5.6 More Analysis
Effect of HQ-Ins Matching. “w/o HQ-Ins Matching” means using detection loss only for bipartite matching [91]. “w/ HQ-Ins Matching” means comprehensively using detection, pose, and segmentation loss for bipartite matching. As shown in Fig. 4, with detection only matching [91], there may be some erroneous cases when one person’s pose is matched to another person. HQ-Ins Matching avoids such errors by comprehensively considering multiple tasks as a whole. More quantitative evaluation can be found in Supplementary.
@@ 6 Conclusion
In this work, we present a unified solution towards single-stage multi-task human-centric perception, called HQNet. The core idea is to learn a unified query representation that encodes local and global appearance features, coarse and fine-grained localization features for each instance. To facilitate model training and evaluation, we introduce a large-scale benchmark, termed COCO-UniHuman benchmark, to unify different representative HCP tasks. We extensively compare our proposed method with several state-of-the-art task-specific and multi-task approaches, and show the effectiveness of our proposed method.
Acknowledgement. This paper is partially supported by the National Key R&D Program of China No.2022ZD0161000 and the General Research Fund of Hong Kong No.17200622 and 17209324.
Appendix 0.A COCO-UniHuman Dataset Statistics
@@
Appendix 0.B COCO-UniHuman Dataset Annotation
@@
0.B.1 Body-based vs face-based annotation strategies.
@@
0.B.2 Two-stage vs one-stage annotation strategies.
@@
Appendix 0.C More Experimental Analysis
@@
0.C.1 General class models vs person-specific models
@@ 0.C.3 Qualitative Results
0.C.4 Attention Visualization
@@
0.C.5 Failure Case Analysis
We have analyzed the main cases where our approach fails in the COCO-UniHuman val set. Fig. A6 shows an overview of common failure cases. In highly crowded and occluded scenes, where people are overlapping, the method tends to miss some targets. Occlusion can also lead to pose estimation errors. There will be high keypoint localization errors on non-typical poses (e.g. upside-down cases). Due to age/gender bias in the dataset, the method may have some erroneous predictions on human attributes. Statues and toys also frequently lead to false positive errors. Most of these issues could be mitigated by adding related data for model training. For example, negative examples could help the network distinguish between humans and other humanoid figures. Adding occluded keypoint annotations could help predict body parts more accurately in occluded scenes.
Figure A6: Common failure cases: (a) missing detection in crowded scenes, (b) false pose detection in occluded scenes, (c) rare pose or appearance, (d) inaccurate or biased age estimation, (e) false positives on statues or toys.
Appendix 0.D More Implementation Details
0.D.1 Loss Functions
@@
0.D.2 Details about Training
@@
0.D.3 Details about Inference
@@
Appendix 0.E Details about Baseline Models
Details about human detection baselines. For human detection, we compare several baseline approaches, i.e. Faster-RCNN [65], IterDETR [100] and DINO [91]. Note that Faster-RCNN and DINO are originally trained to handle general 80 classes (marked with * in Table A3). For fair comparisons, we use MMDetection [9] to re-train and evaluate them on “Person” category using the default experimental setting. Note that MMDetection re-implementation can be a little bit better than the original implementation.
Details about human pose estimation baselines. For human pose estimation, we compare with several representative top-down methods (SBL [82], HRNet [71], Swin [54], ViTPose [86] and PRTR [42]), bottom-up approaches (HrHRNet [11], DEKR [19], and SWAHR [56]) and single-stage approaches (FCPose [57], InsPose [70], PETR [69] and CID [77]). Note that the results of Swin (Swin-L), ViTPose (ViT-L) and CID (R-50-FPN) are from MMPose [13], and other results are from their original papers. Top-down methods generally yield superior performance, but often rely on a separate human detector, incurring redundant computational costs. Specifically, SBL, HRNet, Swin and ViTPose use the same person detector provided by [82], which is a strong Faster-RCNN [65] based detector with detection AP $56.4$ for the “Person” category on the COCO’2017 val set. PRTR applies a DETR-based person detector for human detection, which achieves $50.2$ AP for the whole “Person” category on the COCO’2017 val set. While PRTR introduces an end-to-end variant (E2E-PRTR) optimizing detection and pose jointly, it lags behind separately trained top-down approaches. For pose estimation, the input resolution for SBL, HRNet, and Swin is set as $256\times 192$, while the input resolution for PRTR is $384\times 288$. Bottom-up methods learn instance-agnostic keypoints and then cluster them into corresponding individuals. HrHRNet, DEKR, and SWAHR adopt the strong HRNet-w32 [71] backbone network with an input resolution of $512\times 512$. Single-stage approaches directly predict human body keypoints in a single stage. FCPose, InsPose and PETR adopt R-50 [22] backbone network. The input images are resized to have their shorter sides being 800 and their longer sides less or equal to 1333. For CID, we report both the results of R-50-FPN and HRNet-w32 backbones. The input resolution of CID is $512\times 512$.
@@ Appendix 0.E Details about Baseline Models
Details about mesh estimation baseline. For human mesh estimation, we compare HQNet with state-of-the-art one-stage monocular method ROMP [72], and two-stage method HMR [34] and HMR+ [63]. Following official settings, we train these models on COCO-UniHuman with ResNet50 backbone. As two-stage models require human bbox as the input, in the experiment, we use GT bbox for comparisons.
Appendix 0.F Discussion about Unifying HCP Tasks
0.F.1 General network architecture design
@@ 0.F.3 Co-learning on HCP tasks
Appendix 0.G Discussion about Human Attribute Recognition
@@ 0.G.2 Method
Appendix 0.H Limitations and Future Work
While our work focuses on RGB image based HCP tasks, tasks about video data or multi-modality data (e.g. IR and Depth) also hold significant potential. We encourage future research to explore more comprehensive multi-task human-centric perception.